Weights of 75 pounds and 90 pounds sit at the ends of a 15 foot long board. Where should the fulcrum be placed to balance the board?
I don't know what formula to use to solve this equation.
Any help would be appreciated.
: Weights of 75 pounds and 90 pounds sit at the ends of a 15 foot long board. Where should the fulcrum be placed to balance the board?
I don't know what formula to use to solve this equation.
Any help would be appreciated.

Answer by scott8148(2721)

the weight multiplied by the distance from the fulcrum is a quantity called the "moment"
__ at balance, the moments on either side of the fulcrum are equal

75(x)=90(15-x) __ 75x=1350-90x __ 165x=1350 __ x=90/11

the fulcrum should be 90/11 feet from the 75 lb end